40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 5 13:52:45 2010, Alberto, Update, Computers, Front-ends down. Rebooted 
    Reply  Mon Apr 5 14:10:06 2010, steve, Update, SUS, sus damping restored  
Message ID: 2773     Entry time: Mon Apr 5 14:10:06 2010     In reply to: 2772
Author: steve 
Type: Update 
Category: SUS 
Subject: sus damping restored  

Quote:

This morning, at about 12 Koji found all the front-ends down.

At 1:45pm rebooted ISCEX, ISCEY, SOSVME, SUSVME1, SUSVME2, LSC, ASC, ISCAUX

Then I burtestored ISCEX, ISCEY, ISCAUX to April 2nd, 23:07.

The front-ends are now up and running again.

 I restored damping to all SUSes except ITM-east. The ITMX OSEMs are being used in the clean assembly room.

ELOG V3.1.3-